# Chiral spin liquid in a $\mathbb{Z}_{3}$ Kitaev model

### v2

@@ II $\mathbb{Z}_{3}$ Kitaev Model and symmetries
To see (i), we first note that, since $L_{y}\neq 0\text{ mod }3$, the string operator $\Phi_{2}$, which wraps around the cylinder, is charged under the global $\mathbb{Z}_{3}$ symmetry. To show that the $\mathbb{Z}_{3}$ symmetry is spontaneously broken, it is sufficient to show that $\Phi_{2}$ has long-range correlations. Letting $\Phi_{2,x}$ be the string operator $\Phi_{2}$ supported on sites sharing the same coordinate of the $\mathbf{a}_{1}$-axis (denoted as $x$), we compute the correlator $\langle\Phi_{2,x}\Phi_{2,x+x^{\prime}}^{\dagger}\rangle$. This is non-vanishing in the limit of a large separation between $x$ and $x^{\prime}$, because $\Phi_{2,x}\Phi_{2,x+x^{\prime}}^{\dagger}$ is equivalent to a product of $W_{p}$ operators over the region between $x$ and $x^{\prime}$. Thus, using that the ground states are eigenstates of the $W_{p}$’s, we have $\langle\Phi_{2,x}\Phi_{2,x+x^{\prime}}^{\dagger}\rangle=\omega^{n|x-x^{\prime}|L_{y}}$. A consequence is that, if the global $\mathbb{Z}_{3}$ symmetry is enforced, the system is in a Schrödinger’s cat state.
@@ IV Results on strip geometry
Besides the criticality, we show in Appendix C that there is an additional $3^{L_{x}-1}$-fold degeneracy on the strip geometry, due to non-commuting string-like conserved quantities that extend between the bottom and top of the strip. In the strip geometry, there are also conserved quantities $W_{p,\text{boundary}}$ corresponding to the plaquettes that would connect the top and bottom boundaries if the system were made periodic in the $\mathbf{a}_{2}$-direction. Numerically, we find that the expectation value of $W_{p,\text{boundary}}$ is $\omega$ or $\omega^{2}$ with the same energy (exact or in the infinite $\chi$ limit), for all the strip widths we have considered $L_{y}=2,3,4,5$.
@@ V Discussion and conclusion
    From the EE results on cylinders we find the TEE is $S_{\rm top}=1.28\pm 0.34$.
@@
According to classification of chiral $c=1$ CFTs [45], these theories come in two types: 1) $\mathrm{U}(1)_{2n}$ theories with $n$ a non-zero integer, often referred to as the “circle branch”. 2) Orbifolds of the circle branch, including $\mathbb{Z}_{2}$ orbifolds of $\mathrm{U}(1)_{2n}$ and three exceptional cases (orbifolding non-Abelian groups in SU(2)${}_{1}$). However, due to the bulk-boundary correspondence, the corresponding anyon theory must contain the $\mathbb{Z}_{3}^{(2)}$ theory. Let us consider theories in the circle branch. In order to have $\mathbb{Z}_{3}^{(2)}$ as a subtheory, $n$ must be a multiple of 3, so we write $n=3m$. Then the generator of the $\mathbb{Z}_{3}^{(2)}$ is identified as the $[2m]$ element, with exchange statistics given by $e^{\frac{2\pi im}{3}}$. Therefore, $m$ must be $2$ mod 3, and $n$ must take the form $n=3(3k+2)$ for $k\in\mathbb{Z}$. A similar argument excludes the orbifold theories.
@@
The TEE for the $\mathrm{U}(1)_{6(3k+2)}$ theory is $\ln\sqrt{6\big{|}3k+2\big{|}}$. Compared with the fitted TEE $S_{\rm top}=1.28\pm 0.34$, we find that $k=0$ (the $\mathrm{U}(1)_{12}$ theory) has the closet TEE $\ln\sqrt{12}\approx 1.24$. We note that $k=-2$ (the $\mathrm{U}(1)_{-24}$ theory with TEE $\ln\sqrt{24}\approx 1.58$) is also compatible due to the relatively large errorbar on the TEE. Based on this analysis, we identify $\mathrm{U}(1)_{12}$ as the most likely candidate TO.
@@
Let us discuss the properties of the $\mathrm{U}(1)_{12}=\mathbb{Z}_{3}^{(2)}\boxtimes\mathbb{Z}_{4}^{(3/2)}$ anyon theory. The anyon types can be labeled by an integer mod 12 $[a]$, and $[1]$ is the generator. $[1]$ has exchange statistics $e^{i\pi/12}$. It is easy to check that $[4]$ generates $\mathbb{Z}_{3}^{(2)}$, and $[3]$ generates the $\mathbb{Z}_{4}^{(3/2)}$ part. This theory has 12 degenerate ground states on torus. Starting from one of them, the others can be obtained by applying the corresponding Wilson loop operators. Since the $\mathbb{Z}_{3}^{(2)}$ Wilson loops commute with the Hamiltonian, the ground states are at least three-fold degenerate (as well as all the excited states). The degeneracy associated with $\mathbb{Z}_{4}^{(3/2)}$ is however not expected to be exact and is generally split by quasiparticle tunneling. Indeed, we have managed to find a topologically quasi-degenerate ground state on a width-3 cylinder (discussed in Appendix E). Nevertheless, due to the high topological degeneracy, finding all topologically degenerate ground states seem to be a daunting task.
@@ VI Acknowledgements
We acknowledge conversations with Johannes Hauschild, Xiao-Yu Dong, Hui-Ke Jin and Hong-Hao Tu. The MPS calculations were performed using the TeNPy Library (version 0.9.0) [48]. L.M.C. and P.Y. were supported by NSFC Grant No. 12074438, Guangdong Basic and Applied Basic Research Foundation under Grant No. 2020B1515120100, and the Open Project of Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No. 2022B1212010008. M.C. acknowledges support from NSF under award number DMR-1846109. J.-Y.C. was supported by Open Research Fund Program of the State Key Laboratory of Low-Dimensional Quantum Physics (Project No. KF202207), Fundamental Research Funds for the Central Universities, Sun Yat-sen University (Project No. 23qnpy60), a startup fund from Sun Yat-sen University, the Innovation Program for Quantum Science and Technology 2021ZD0302100, and National Natural Science Foundation of China (NSFC) (Grant No. 12304186). This work was also supported by the Fundamental Research Funds for Central Universities (22qntd3005). The Calculations reported were performed on resources provided by the Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices, No. 2022B1212010008.
@@ Appendix C Degeneracy on the strip geometry
Figure 7: The conserved quantities of the $\mathbb{Z}_{3}$ Kitaev model on a strip geometry are defined in terms of the operators $W_{i}$ and $\tilde{W}_{i}$, where $i$ indexes the bivalent vertices at the top of the strip. The operator $W_{i}$, shown in red, is a product of $T^{y}$ and $(T^{y})^{\dagger}$ operators, while $\tilde{W}_{i}$, shown in blue, is a product of $T^{x}$, $T^{y}$, and $T^{z}$.
Here, we argue that there is at least a $3^{L_{x}-1}$-fold degeneracy of the $\mathbb{Z}_{3}$ Kitaev model on a strip-like geometry with length $L_{x}$ in the ${\mathbf{a}_{1}}$-direction. In particular, we identify $2(L_{x}-1)$ conserved quantities of the Hamiltonian, which come in non-commuting pairs. Their commutation relations imply that there is a 3-fold degeneracy enforced by each pair.
To specify the conserved quantities, we begin by indexing the bivalent vertices at the top of the strip from 1 to $L_{x}$, as shown in Fig. 7. For $i\in\{1,\ldots,L_{x}-1\}$, we define $W_{i}$ and $\tilde{W}_{i}$ to be the string operators ending at the vertex $i$, as depicted in Fig. 7. We then define the following two conserved quantities, for each $i$ in $\{1,\ldots,L_{x}-1\}$:
$\bar{T}_{i}^{z}\equiv(W_{i})^{i},\quad\bar{T}_{i}^{x}\equiv\prod_{j=1}^{i}(\tilde{W}_{j})^{j}.$ (8)
Notably, these operators commute with all of the Hamiltonian terms and satisfy the same relations as the operators $T^{z}$ and $T^{x}$:
$(\bar{T}_{i}^{x})^{3}=(\bar{T}_{i}^{z})^{3}=1,\quad\bar{T}_{i}^{z}\bar{T}_{j}^{x}=\begin{cases}\omega\bar{T}_{j}^{x}\bar{T}_{i}^{z}&i=j\\ \bar{T}_{j}^{x}\bar{T}_{i}^{z}&i\neq j.\end{cases}$ (9)
The commutation relations of the conserved quantities $\bar{T}_{i}^{z}$ and $\bar{T}_{j}^{x}$ implies that there is a 3-fold degeneracy for each $i\in\{1,\ldots,L_{x}-1\}$. Therefore, the $\mathbb{Z}_{3}$ Kitaev model has at least a $3^{L_{x}-1}$-fold degeneracy on a strip. We note that this degeneracy can be lifted by including single site terms in the Hamiltonian for each bivalent vertex at the boundary.
@@ Appendix D Additional data for $L_{y}=5$ strip
Here we show additional evidence that the critical behavior can also be observed on a $L_{y}=5$ strip. In Fig. 8, we compare the maximal correlation and entanglement entropy of a $L_{y}=5$ strip and cylinder. Both correlation length and entanglement entropy grows much quicker (also with larger values) on the strip geometry, without any sign of saturation. This scenario is consistent with the data for $L_{y}=3,4$ shown in the main manuscript.
@@ Appendix E Quasi-degeneracy on $L_{y}=3$ cylinder
A comparison of $|\Psi_{\mathrm{I}}\rangle$ and $|\Psi_{\epsilon}\rangle$ is shown in Fig. 9. We identify $|\Psi_{\epsilon}\rangle$ as an approximately degenerate ground state on the finite-circumference cylinder. Because of the finite width, the topological degeneracy would be split by virtual quasiparticle tunneling effects, which explains the energy difference as shown in the inset of Fig. 9. Moreover, on YC3333 cylinder, the maximal correlation length of $|\Psi_{\epsilon}\rangle$ is converging with increasing bond dimension $\chi$ (shown in Fig. 9 inset), implying that the $|\Psi_{\epsilon}\rangle$ is also gapped, although the value is larger than that of $|\Psi_{\mathrm{I}}\rangle$. On the other hand, $|\Psi_{\epsilon}\rangle$ has a higher half cylinder entanglement entropy than $|\Psi_{\mathrm{I}}\rangle$, with the difference approaching 0.36 with increasing $\chi$. However, we should caution that we have not been able to find similar approximately degenerate states on wider cylinders.

### v3

@@ II $\mathbb{Z}_{3}$ Kitaev Model and symmetries
To see (i), we first note that, since $L_{y}\neq 0\text{ mod }3$, the string operator $\Phi_{2}$, which wraps around the cylinder, is charged under the global $\mathbb{Z}_{3}$ symmetry. To show that the $\mathbb{Z}_{3}$ symmetry is spontaneously broken, it is sufficient to show that $\Phi_{2}$ has long-range correlations. Letting $\Phi_{2,x}$ be the string operator $\Phi_{2}$ supported on sites sharing the same coordinate of the $\mathbf{a}_{1}$-axis (denoted as $x$), we compute the correlator $\langle\Phi_{2,x}\Phi_{2,x^{\prime}}^{\dagger}\rangle$. This is non-vanishing in the limit of a large separation between $x$ and $x^{\prime}$, because $\Phi_{2,x}\Phi_{2,x^{\prime}}^{\dagger}$ is equivalent to a product of $W_{p}$ operators over the region between $x$ and $x^{\prime}$. Thus, using that the ground states are eigenstates of the $W_{p}$’s, we have $\langle\Phi_{2,x}\Phi_{2,x^{\prime}}^{\dagger}\rangle=\omega^{n|x-x^{\prime}|L_{y}}$. A consequence is that, if the global $\mathbb{Z}_{3}$ symmetry is enforced, the system is in a Schrödinger’s cat state.
@@ IV Results on strip geometry
Besides the criticality, we show in Appendix C that there is an additional $3^{L_{x}}$-fold degeneracy on the strip geometry, due to non-commuting string-like conserved quantities that extend between the bottom and top of the strip. In the strip geometry, there are also conserved quantities $W_{p,\text{boundary}}$ corresponding to the plaquettes that would connect the top and bottom boundaries if the system were made periodic in the $\mathbf{a}_{2}$-direction. Numerically, we find that the expectation value of $W_{p,\text{boundary}}$ is $\omega$ or $\omega^{2}$ with the same energy (exact or in the infinite $\chi$ limit), for all the strip widths we have considered $L_{y}=2,3,4,5$.
@@ V Discussion and conclusion
    From the entanglement entropy results on cylinders we find the TEE is $S_{\rm top}=1.28\pm 0.34$.
@@
According to classification of chiral $c=1$ CFTs [45], these theories come in two types: 1) $\mathrm{U}(1)_{2n}$ theories with $n$ a non-zero integer, often referred to as the “circle branch”. 2) Orbifolds of the circle branch, including $\mathbb{Z}_{2}$ orbifolds of $\mathrm{U}(1)_{2n}$ and three exceptional cases (orbifolding non-Abelian groups in SU(2)1). We will focus on the circle branch, which will turn out to be more relevant for this system. The $\mathrm{U}(1)_{2n}$ CFT can be described by a chiral compact boson. They can arise as the edge theory of a 2+1D chiral topological phase, the bulk of which is described by the $\mathrm{U}(1)_{2n}$ anyon theory. An example is the $\nu=\frac{1}{2}$ bosonic Laughlin state (also known as the Kalmeyer-Laughlin state) with $\mathrm{U}(1)_{2}$ chiral edge CFT. Let us enumerate the anyon content of the $\mathrm{U}(1)_{2n}$ theory. The different anyon types can be labeled by an integer defined mod $2n$, denoted as $[a]$ where $a\in\mathbb{Z}$, with $[0]$ representing the trivial excitations. $[a]$ has exchange statistics $e^{\frac{i\pi a^{2}}{2n}}$. The fusion rule of anyons is given by addition, so $[a]\times[b]=[a+b]$. The anyon types form a $\mathbb{Z}_{2n}$ group under fusion, with $[1]$ as the generator. Using the notations in [24], the theory is denoted as $\mathbb{Z}_{2n}^{(1/2)}$.
However, we have also established through the exact 1-form symmetry of the lattice model that the bulk anyon theory must contain the $\mathbb{Z}_{3}^{(2)}$ as a subtheory, which places strong constraint on the value of $n$. First, to have $\mathbb{Z}_{3}^{(2)}$ as a subtheory, $\mathbb{Z}_{3}$ must be a subgroup of $\mathbb{Z}_{2n}$, meaning that $n$ must be a multiple of $3$, so we write $n=3m$. The unique $\mathbb{Z}_{3}$ subgroup in $\mathbb{Z}_{6m}$ is $\{[0],[2m],[4m]\}$. The self statistics of the generator $[2m]$ is $e^{\frac{i\pi}{6m}(2m)^{2}}=e^{\frac{2\pi i}{3}m}$, therefore to match $\mathbb{Z}_{3}^{(2)}$, $m$ must be 2 mod 3, and $n$ must take the form $n=3(3k+2)$ for $k\in\mathbb{Z}$. We note that a similar analysis can exclude the orbifold branch.
@@
To further constrain the possible values of $k$, we consider the TEE. It is known that the TEE for the $\mathrm{U}(1)_{6(3k+2)}$ theory is $\ln\sqrt{6\big{|}3k+2\big{|}}$. Compared with the fitted TEE $S_{\rm top}=1.28\pm 0.34$, we find that $k=0$ (the $\mathrm{U}(1)_{12}$ theory) has the closet TEE $\ln\sqrt{12}\approx 1.24$. We note that $k=-2$ (the $\mathrm{U}(1)_{-24}$ theory with TEE $\ln\sqrt{24}\approx 1.58$) is also compatible, due to the relatively large errorbar on the numerically computed TEE. While $\mathrm{U}(1)_{-24}$ can not be excluded based on our data so far, we believe the fact that $\mathrm{U}(1)_{-24}$ has many more anyons and the TEE is close to the upper bound set by the errorbar makes it less likely to be the the actual ground state topological order. Based on this analysis, we identify $\mathrm{U}(1)_{12}$ as the most likely candidate topological order.
@@
Let us discuss the properties of the $\mathrm{U}(1)_{12}=\mathbb{Z}_{3}^{(2)}\boxtimes\mathbb{Z}_{4}^{(3/2)}$ anyon theory. The anyon content has been discussed above. It is easy to check that $[4]$ generates $\mathbb{Z}_{3}^{(2)}$, and $[3]$ generates the $\mathbb{Z}_{4}^{(3/2)}$ part. This theory has 12 degenerate ground states on torus. Starting from one of them, the others can be obtained by applying the corresponding Wilson loop operators. Since the $\mathbb{Z}_{3}^{(2)}$ Wilson loops commute with the Hamiltonian, the ground states are at least three-fold degenerate (as well as all the excited states). The degeneracy associated with $\mathbb{Z}_{4}^{(3/2)}$ is however not expected to be exact and is generally split by quasiparticle tunneling. Indeed, we have managed to find a topologically quasi-degenerate ground state on a width-3 cylinder (discussed in Appendix E). Nevertheless, due to the high topological degeneracy, finding all topologically degenerate ground states seems to be a daunting task.
@@ VI Acknowledgements
We acknowledge conversations with Johannes Hauschild, Xiao-Yu Dong, Hui-Ke Jin and Hong-Hao Tu. The MPS calculations were performed using the TeNPy Library (version 0.9.0) [48]. L.M.C. and P.Y. were supported by NSFC Grant No. 12074438, Guangdong Basic and Applied Basic Research Foundation under Grant No. 2020B1515120100, and the Open Project of Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices under Grant No. 2022B1212010008. M.C. acknowledges support from NSF under award number DMR-1846109. J.-Y.C. was supported by Open Research Fund Program of the State Key Laboratory of Low-Dimensional Quantum Physics (Project No. KF202207), Fundamental Research Funds for the Central Universities, Sun Yat-sen University (Project No. 23qnpy60), a startup fund from Sun Yat-sen University, the Innovation Program for Quantum Science and Technology 2021ZD0302100, Guangzhou Basic and Applied Basic Research Foundation (grant No. 2024A04J4264), and National Natural Science Foundation of China (NSFC) (Grant No. 12304186). This work was also supported by the Fundamental Research Funds for Central Universities (22qntd3005). The Calculations reported were performed on resources provided by the Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices, No. 2022B1212010008.
@@ Appendix C Degeneracy on the strip geometry
Figure 7: The conserved quantities of the $\mathbb{Z}_{3}$ Kitaev model on a strip geometry are generated by the operators $W_{i}$ and $\tilde{W}_{i}$, where $i$ indexes the bivalent vertices at the top of the strip. The operator $W_{i}$, shown in red, is a product of $T^{y}$ and $(T^{y})^{\dagger}$ operators, while $\tilde{W}_{i}$, shown in blue, is a product of $T^{x}$, $T^{y}$, and $T^{z}$. If $L_{x}$ is even, then each pair of $W_{i}$ and $\tilde{W}_{i}$, for even $i$, generates the same operator algebra as a pair of $T^{x}$ and $T^{y}$ operators. This implies that each pair enforces a 3-fold degeneracy of the ground state. Note that $\tilde{W}_{i}$ and $\tilde{W}_{i-1}$ are non-commuting, so to ensure that the pairs are independent, we consider only those at even sites. The subsystem code calculation suggests that it is possible to find $L_{x}$ independent non-commuting pairs.
Here, we argue that there is at least a $3^{L_{x}}$-fold degeneracy of the $\mathbb{Z}_{3}$ Kitaev model on a strip-like geometry with length $L_{x}$ in the ${\mathbf{a}_{1}}$-direction and width $L_{y}$ in the ${\mathbf{a}_{2}}$-direction. To see this, we view the $\mathbb{Z}_{3}$ Kitaev model as a subsystem code whose gauge group is generated by products of the Hamiltonian terms. For an introduction to subsystem codes and a more careful description of the code associated to the $\mathbb{Z}_{3}$ Kitaev model, we refer to Ref. [23].
More specifically, we show that the subsystem code has a logical subsystem of dimension $3^{L_{x}}$. This implies that there are $L_{x}$ non-commuting pairs of conserved quantities of the $\mathbb{Z}_{3}$ Kitaev model that satisfy the same commutation relations as $T^{x}$ and $T^{z}$. Consequently, the model must have a $3^{L_{x}}$-fold degeneracy. To support the calculation that follows, we identify $L_{x}/2$ explicit pairs of non-commuting conserved quantities in Fig. 7, for a strip of even length $L_{x}$. This suffices to show that there is at least a $3^{L_{x}/2}$-fold degeneracy.
We now compute the dimension of the logical subsystem of the subsystem code by comparing the number of physical qutrits (i.e., three-level local degrees of freedom) to the number of stabilized qutrits and gauge qutrits. There are two qutrits for every plaquette. Due to the open boundary conditions, there are an additional $2L_{x}$ qutrits associated to the edges of the strip. In total, the number of qutrits $\mathbf{N}_{Q}$ is $\mathbf{N}_{Q}=2L_{x}L_{y}$.
Next, we count the number of stabilized qutrits and the number of gauge qutrits. The only stabilizers are the conserved quantities $W_{p}$ for every plaquette, each of which stabilizes a single qutrit. There are $L_{x}(L_{y}-1)$ plaquettes, so the number of stabilized qutrits $\mathbf{N}_{S}$ is $\mathbf{N}_{S}=L_{x}(L_{y}-1)$. There are three gauge generators for every plaquette. There are an additional two for every plaquette on an edge of the strip. Therefore, the number of gauge qutrits $\mathbf{N}_{G}$ is
$\mathbf{N}_{G}=[(3L_{x}(L_{y}-1)+2L_{x})-\mathbf{N}_{S}]/2=L_{x}L_{y}.$ (8)
Here, we have subtracted $\mathbf{N}_{S}$, since the stabilizers are generated by the gauge generators, and we have divided by two to account for the fact that the operator algebra of a qutrit has two generators $T^{x}$ and $T^{z}$.
The number of logical qutrits $\mathbf{N}_{L}$ is then given by the formula
$\mathbf{N}_{L}=\mathbf{N}_{Q}-\mathbf{N}_{S}-\mathbf{N}_{G}=L_{x}.$ (9)
Thus, there are $L_{x}$ logical qutrits, which means that there are $2L_{x}$ generators of the logical group; each of which is a conserved quantity of the $\mathbb{Z}_{3}$ Kitaev model. Therefore, we conclude that there is a $3^{L_{x}}$-fold degeneracy on a strip geometry.
We note that this degeneracy can be lifted by including a single site term in the Hamiltonian for each bivalent vertex at the boundary.
@@ Appendix D Additional data for $L_{y}=5$ strip
Here we show additional evidence that the critical behavior can also be observed on a $L_{y}=5$ strip. In Fig. 8, we compare the maximal correlation length and entanglement entropy of a $L_{y}=5$ strip and cylinder. Both correlation length and entanglement entropy grows much quicker (also with larger values) on the strip geometry, without any sign of saturation. This scenario is consistent with the data for $L_{y}=3,4$ shown in the main manuscript.
@@ Appendix E Quasi-degeneracy on $L_{y}=3$ cylinder
A comparison of $|\Psi_{\mathrm{I}}\rangle$ and $|\Psi_{\epsilon}\rangle$ is shown in Fig. 9. We identify $|\Psi_{\epsilon}\rangle$ as an approximately degenerate ground state on the finite-circumference cylinder. Because of the finite width, the topological degeneracy would be split by virtual quasiparticle tunneling effects, which explains the energy difference as shown in the inset of Fig. 9. Moreover, on YC3333 cylinder, the maximal correlation length of $|\Psi_{\epsilon}\rangle$ is converging with increasing bond dimension $\chi$ (shown in Fig. 9 inset), implying that the $|\Psi_{\epsilon}\rangle$ is also gapped, although the value is larger than that of $|\Psi_{\mathrm{I}}\rangle$. On the other hand, $|\Psi_{\epsilon}\rangle$ has a higher half cylinder entanglement entropy than $|\Psi_{\mathrm{I}}\rangle$, with the difference approaching 0.36 with increasing $\chi$.